OpenStudy (anonymous):

f(x,y) = 2xy -x -y +1 Find the inverse. The answer is: g(x) = x/(2x-1) but I can't quite figure out why. Can anyone point me in the right direction here please.

OpenStudy (amistre64):

the inverse of a multivariable function? or is this spose to represent something in the xy plane?

OpenStudy (amistre64):

f(x,y) = 2xy -x -y +1 with respect to x, maybe z = 2xy -x -y +1 z = 2kx -x -k +1 z = (2k-1)x -k +1 z + k - 1= (2k-1)x (z + k - 1)/(2k-1 )= x (z + y - 1)/(2y-1 )= x

OpenStudy (amistre64):

with respect to y would amount to the same results, just swapping x for y

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=inverse+2xy+-x+-y+%2B1+ seems to be something to it ...